OpenStudy (anonymous):

8(t+2)-3(t-4)=6(t-7)+8

jimthompson5910 (jim_thompson5910):

8(t+2)-3(t-4)=6(t-7)+8 8t+16-3t+12=6t-42+8 28+5t=6t-42+8 28+5t=-34+6t 5t=-34+6t-28 5t-6t=-34-28 -t=-34-28 -t=-62 t=-62/(-1) t=62

OpenStudy (across):

Jim is correct. \[8(t+2)-3(t-4)=6(t-7)+8\]\[8t+16-3t+12=6t-42+8\]\[5t+28=6t-34\]\[62=t\]